OpenStudy (anonymous):

What is the value of this summation? 4 Σ (2i+6i^2) I=1 I

OpenStudy (anonymous):

A: 180 B: 200 C: 220 D: 240

OpenStudy (anonymous):

\[\sum_{i=1}^{4}\left( 2i+6i ^{2} \right)=2\left( 1+2+3+4 \right)+6\left( 1^{2}+2^{2}+3^{2}+4^{2} \right)\] \[=2\left( 10 \right)+6\left( 1+4+9+16 \right)=20+6*30=200\]

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

yw